Problem Solving: Solve the following problems. Show your complete solution and box your final answer.


A wall clock is placed on a wall so it can easily be seen. The gravitational potential energy experienced by the wall clock is 2.94 joules. If the mass of the wall clock is 200g, how high is


the clock relative to the ground?


1
Expert's answer
2021-11-29T11:40:38-0500
"PE=mgh,""h=\dfrac{PE}{mg}=\dfrac{2.94\ J}{0.2\ kg\times9.8\ \dfrac{m}{s^2}}=1.5\ m."